A cylindrical jug that carries 5 ℓ of water when it is filled to the brim, has a height of 20 cm


1
Expert's answer
2022-05-10T08:16:26-0400
"V=\dfrac{\pi d^2}{4}h"

"d=2\sqrt{\dfrac{V}{\pi h}}"

Given "V=5000{cm}^3, h=20cm"


"d=2\sqrt{\dfrac{5000{cm}^3}{\pi (20cm)}}=10\sqrt{\dfrac{10}{\pi}}cm\approx17.841cm"